# TFB: Towards Comprehensive and Fair Benchmarking of Time Series Forecasting Methods

### v2

@@ Abstract.
Time series are generated in diverse domains such as economic, traffic, health, and energy, where forecasting of future values has numerous important applications. Not surprisingly, many forecasting methods are being proposed. To ensure progress, it is essential to be able to study and compare such methods empirically in a comprehensive and reliable manner. To achieve this, we propose TFB, an automated benchmark for Time Series Forecasting (TSF) methods. TFB advances the state-of-the-art by addressing shortcomings related to datasets, comparison methods, and evaluation pipelines: 1) insufficient coverage of data domains, 2) stereotype bias against traditional methods, and 3) inconsistent and inflexible pipelines. To achieve better domain coverage, we include datasets from 10 different domains: traffic, electricity, energy, the environment, nature, economic, stock markets, banking, health, and the web. We also provide a time series characterization to ensure that the selected datasets are comprehensive. To remove biases against some methods, we include a diverse range of methods, including statistical learning, machine learning, and deep learning methods, and we also support a variety of evaluation strategies and metrics to ensure a more comprehensive evaluations of different methods. To support the integration of different methods into the benchmark and enable fair comparisons, TFB features a flexible and scalable pipeline that eliminates biases.
Next, we employ TFB to perform a thorough evaluation of 21 Univariate Time Series Forecasting (UTSF) methods on 8,068 univariate time series and 14 Multivariate Time Series Forecasting (MTSF) methods on 25 datasets. The results offer a deeper understanding of the forecasting methods, allowing us to better select the ones that are most suitable for particular datasets and settings. Overall, TFB and this evaluation provides researchers with improved means of designing new TSF methods. The benchmark code and data are available at https://github.com/decisionintelligence/TFB.
Corresponding Author: Jilin Hu (jlhu@dase.ecnu.edu.cn)
Reference Format:
Xiangfei Qiu, Jilin Hu, Lekui Zhou, Xingjian Wu, Junyang Du, Buang Zhang, Chenjuan Guo, Aoying Zhou, Christian S. Jensen, Zhenli Sheng and Bin Yang.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license.
@@ 1. Introduction
As part of the ongoing digitalization, time series are generated in a variety of domains, such as economic (Sezer et al., 2020; Huang et al., 2022), traffic (Xu et al., 2023; Pedersen et al., 2020; Hu et al., 2018; Mo et al., 2022; Wan et al., 2022; Yang et al., 2023; Hu et al., 2019; Guo et al., 2020; Hu et al., 2017), health (Wei et al., 2022; Miao et al., 2021; Lee et al., 2017; Tran et al., 2019), energy (Alvarez et al., 2010; Guo et al., 2015), and AIOps (Wang et al., 2024; Campos et al., 2022; Wang et al., 2023; Qi et al., 2023; Kieu et al., 2022a, b; Zhao et al., 2022). Time Series Forecasting (TSF) is essential in key applications in these domains (Pan et al., 2023; Guo et al., 2014; Yu et al., 2023; Yang et al., 2021). Given historical observations, it is valuable if we can know the future values ahead of time. Correspondingly, TSF has been firmly established as an active research field, witnessing the proposal of numerous methods.
Time series organize data points chronologically and are either univariate or multivariate depending on the number of variables in each data point. Accordingly, TSF methods can be classified as either Univariate Time Series Forecasting (UTSF) or Multivariate Time Series Forecasting (MTSF) methods. Among early methods, Autoregressive Integrated Moving Average (ARIMA) (Box and Pierce, 1970) and Vector Autoregression (VAR) (Toda and Phillips, 1994) are arguably the most popular univariate and multivariate forecasting methods, respectively. Subsequent methods that exploit machine learning, e.g., XGBoost (Chen and Guestrin, 2016; Zhang et al., 2021) and Random Forest (Breiman, 2001; Mei et al., 2014) offer better performance than the early methods. Most recently, methods based on deep learning have demonstrated state-of-the-art (SOTA) forecasting performance on a variety of datasets (Zhou et al., 2021; Nie et al., 2022; Wu et al., 2022; Zhang and Yan, 2022; Chen et al., 2024; Yao et al., 2023; Wu et al., 2021b, 2023; Zhao et al., 2023; Cheng et al., 2023; Cirstea et al., 2022b, 2019, 2021; Miao et al., 2024).
@@
Figure 1. Visualization of data with different characteristics.
@@
Issue 1. Insufficient Coverage of Data Domains. Time series from different domains may exhibit diverse characteristics. Figure 1a depicts a time series from the environment domain called AQShunyi (Zhang et al., 2017) that records temperature information at hourly intervals, exhibiting a distinct seasonal pattern. This pattern is reasonable in this scenario because temperatures in nature often cycle around the year. Figure 1b shows a time series from FRED-MD (McCracken and Ng, 2016) belongs to economic domain that describes the monthly macroeconomic from 114 regional, national, and international sources with a clear increasing tendency. This may attributed to overall economic stability with minimal fluctuations, reflecting sustained growth in the macroeconomic indicators. Figure 1c depicts a series among Electricity (Trindade, 2015) which comes from electricity domain and has a significant change in the data at a certain point in time, which might indicate an abrupt event, etc. However, these simple patterns are only the tip of the iceberg, and time series from different domains may exhibit much more complex patterns that either combine the above characteristics or are entirely different. Therefore, using only limited domains results in limited coverage of time series characteristics, which cannot offer a full picture.
@@
Datasets
VAR
LR
PatchTST
NLinear
FEDformer
Crossformer
NASDAQ 0.462
0.705
0.567
0.522
0.547
0.745
Wind
0.620
0.583
0.652
0.640
0.697
0.590
ILI
1.012
14.244
0.835 0.919
1.020
1.096
@@
Table 2. Impact of batch sizes with “drop last.” Size PatchTST DLinear FEDformer 1 0.4203 0.4874 0.4120 32 0.4138 0.4831 0.4084 64 0.3999 0.4726 0.4022 128 0.3750 0.4539 0.3921 256 0.3561 0.4360 0.3825 512 0.3483 0.4251 0.3736
Figure 4. “Drop last” illustration.
Issue 3. Lack of consistent and flexible pipelines. The performance of different methods varies with changing experimental settings, e.g., splits among training/validation/testing data, the choice of normalization method, and the selection of hyper-parameter settings. For example, implementations of existing methods often employ a “Drop Last” trick in the testing phase (Zhou et al., 2022b; Nie et al., 2022; Wang et al., 2022). To accelerate the testing, it is common to split the data into batches. However, if we discard the last incomplete batch with fewer instances than the batch size, this causes unfair comparisons. For example, in Figure 4, the ETTh2 (Zhou et al., 2021) has a testing series of length 2,880, and we need to predict 336 future time steps using a look-back window of size 512. If we select the batch sizes to be 32, 64, and 128, the number of samples in the last batch are 17, 49, and 113, respectively. Discarding those last-batch testing samples is inappropriate unless all methods use the same strategy. Table 4 shows the testing results of PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023) and FEDformer (Zhou et al., 2022b) on the ETTh2 with different batch sizes and the “Drop Last” trick turned on. We observe that the performance of the methods changes when varying the batch size. However, some existing benchmarks, such as TSlib (Wu et al., 2022), adopt the “Drop Last.” TSlib uses almost the same batch size across all methods in the evaluation phase to achive consistent comparisons. However, when the users select different batch sizes inadvertently, inconsistent comparisons occur. In addition, the pipelines in most benchmarks are inflexible and do not support simultaneous evaluation of statistical learning, machine learning, and deep learning methods. We argue that it is crucial to ensure a consistent and flexible pipeline so that methods are evaluated in the same setting, thereby improving the fairness of the findings.
Robust and extensive benchmarks enable researchers to evaluate proposals for new methods more rigorously across diverse range of datasets, which is crucial for advancing the state-of-the-art (Tan et al., 2020). For example, ImageNet (Deng et al., 2009), that encompasses a substantial dataset has been instrumental in ensuring progress in computer vision. ImageNet has established itself as a standard for assessing methods in image processing due to its support for rigorous evaluation. Table 3 compares existing benchmarks for TSF according to seven properties. No existing benchmark possesses all properties.
@@
Property Benchmark M3 (Makridakis and Hibon, 2000) M4 (Makridakis et al., 2018) LTSF-Linear (Zeng et al., 2023) TSlib (Wu et al., 2022) BasicTS (Liang et al., 2022) BasicTS+ (Shao et al., 2023) Monash (Godahewa et al., 2021) Libra (Bauer et al., 2021) TFB (Ours)
Univariate forecasting $\surd$ $\surd$
@@
$\surd$ $\surd$ $\surd$
Multivariate forecasting
@@
$\times$
$\surd$ $\surd$ $\surd$ $\surd$
$\times$
$\times$
$\surd$
Statistical method $\surd$ $\surd$
$\times$
$\times$
$\times$
$\times$
$\surd$ $\surd$ $\surd$
Machine learning method $\surd$ $\surd$
@@
$\surd$ $\surd$ $\surd$
Deep learning method
$\times$
$\surd$ $\surd$ $\surd$ $\surd$ $\surd$
$\times$
@@
$\times$
$\surd$
Flexible & scalable pipeline
@@ 3. Preliminaries
(4) $\vspace{-1mm}\footnotesize F=\mathit{Catch22}(X)$
@@
(5) $\vspace{-1mm}\footnotesize P=\left\{r(F^{i},F^{j})\mid 1\leq i\leq N,i+1\leq j\leq N,i,j\in N^{\ast}\right\}$
@@ 4.1.1. Dataset overview
Univariate time series. The univariate datasets are carefully curated from 16 open-source datasets, thus covering dozens of domains. To fully reflect the complexity of real-world time series, we employed Pattern Frequency Analysis (PFA) (Lu et al., 2007), a variation of Principal Component Analysis (PCA) (Bro and Smilde, 2014). PFA preserves the original values of individual time series data points. We employ the concept of explained variance, representing the ratio between the variance of a single time series and the sum of variances across all individual time series. A threshold $t$ for the explained variance is set to 0.9. This implies that for each dataset, we choose to retain the minimum number of time series required to encompass 90% of the variance contributed by the remaining time series. As a result, the selected data exhibits pronounced heterogeneity. Compared to datasets with strong homogeneity, it can better reflect methods performance. In the end, we select 8,068 time series to enable the combined dataset to capture the diversity of real-world time series. Statistical information is reported in Table 4.
@@ 4.2. Comparison Methods
To investigate the strengths and limitations of different forecasting methods, we include 22 methods that can be categorized as statistical learning, machine learning, and deep learning methods. In terms of statistical learning methods, we include ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Kalman Filter (KF) (Harvey, 1990), and VAR (Toda and Phillips, 1994). Among the machine learning methods, we include XGBModel (XGB) (Herzen et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), and Random Forest (RF) (Breiman, 2001). Finally, we further split the deep learning methods into RNN-based models (RNN (Kim et al., 2021)), CNN-based models (MICN (Wang et al., 2022), TimesNet (Wu et al., 2022), and TCN (Bai et al., 2018)), MLP-based models (NLinear (Zeng et al., 2023), DLinear (Zeng et al., 2023), TiDE (Das et al., 2023), N-HiTS (Challu et al., 2023), and N-BEATS (Oreshkin et al., 2019)), Transformer-based models (PatchTST (Nie et al., 2022), Crossformer (Zhang and Yan, 2022), and FEDformer (Zhou et al., 2022b), Non-stationary Transformer (Stationary) (Liu et al., 2022), Informer (Zhou et al., 2021), and Triformer (Cirstea et al., 2022a)), and Model-Agnostic models (FiLM (Zhou et al., 2022a)). Figure 6 gives an overview of the diverse technical approaches of the methods covered by the benchmark. This wide selection of comparison methods aims to offer a deeper understanding of the technical nuances and performance variations across methods.
@@ 4.3.1. Evaluation strategies
Rolling forecasting. In Rolling forecasting, illustrated in Figure 6(b), the blue squares represent historical data, the green squares represent the forecasting horizon, and the white squares represent the remaining data in the time series. During rolling forecast, excluding the last iteration, the historical data expands by a fixed number of steps (called the stride) in each iteration. In the final iteration, the historical data is extended to cover the entire time series along with the forecasting horizon. In each iteration of the inference process, the forecasting model is applied based on the historical data to forecast the designated forecasting horizon. Then, we calculat the average of the error metrics for each iteration. In statistical learning methods (e.g., ARIMA (Box and Pierce, 1970), ETS (Federico Garza, 2022)), it is common to use the entire or a subset of the historical data for training and then make forecasting during each iteration. In contrast, in deep learning or machine learning methods, each iteration often involves using only the last portion of the historical data, with a length equal to the look-back windows size, for inference to make forecasting. The current evaluation strategy for TSF, such as TimesNet (Wu et al., 2022), aligns with our defined criteria.

### v3

@@ Abstract.
Time series are generated in diverse domains such as economic, traffic, health, and energy, where forecasting of future values has numerous important applications. Not surprisingly, many forecasting methods are being proposed. To ensure progress, it is essential to be able to study and compare such methods empirically in a comprehensive and reliable manner. To achieve this, we propose TFB, an automated benchmark for Time Series Forecasting (TSF) methods. TFB advances the state-of-the-art by addressing shortcomings related to datasets, comparison methods, and evaluation pipelines: 1) insufficient coverage of data domains, 2) stereotype bias against traditional methods, and 3) inconsistent and inflexible pipelines. To achieve better domain coverage, we include datasets from 10 different domains : traffic, electricity, energy, the environment, nature, economic, stock markets, banking, health, and the web. We also provide a time series characterization to ensure that the selected datasets are comprehensive. To remove biases against some methods, we include a diverse range of methods, including statistical learning, machine learning, and deep learning methods, and we also support a variety of evaluation strategies and metrics to ensure a more comprehensive evaluations of different methods. To support the integration of different methods into the benchmark and enable fair comparisons, TFB features a flexible and scalable pipeline that eliminates biases. Next, we employ TFB to perform a thorough evaluation of 21 Univariate Time Series Forecasting (UTSF) methods on 8,068 univariate time series and 14 Multivariate Time Series Forecasting (MTSF) methods on 25 datasets. The results offer a deeper understanding of the forecasting methods, allowing us to better select the ones that are most suitable for particular datasets and settings. Overall, TFB and this evaluation provide researchers with improved means of designing new TSF methods.
This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 17, No. 9 ISSN 2150-8097.
doi:10.14778/3665844.3665863
PVLDB Reference Format:
Xiangfei Qiu, Jilin Hu, Lekui Zhou, Xingjian Wu, Junyang Du, Buang Zhang, Chenjuan Guo, Aoying Zhou, Christian S. Jensen, Zhenli Sheng and Bin Yang. PVLDB, 17(9): 2363 - 2377, 2024.
doi:10.14778/3665844.3665863
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/decisionintelligence/TFB.
@@ 1. Introduction
As part of the ongoing digitalization, time series are generated in a variety of domains, such as economic (Sezer et al., 2020; Huang et al., 2022), traffic (Xu et al., 2023; Lin et al., 2023b; Hu et al., 2018; Mo et al., 2022; Lin et al., 2024; Sun et al., 2023; Lin et al., 2021; Wan et al., 2022; Yang et al., 2023; Hu et al., 2019; Guo et al., 2020; Hu et al., 2017), health (Wei et al., 2022; Miao et al., 2021; Lee et al., 2017; Tran et al., 2019), energy (Alvarez et al., 2010; Guo et al., 2015), and AIOps (Campos et al., 2022; Wang et al., 2023; Qi et al., 2023; Campos et al., 2024; Kieu et al., 2022; Zhao et al., 2022). Time Series Forecasting (TSF) is essential in key applications in these domains (Pan et al., 2023; Guo et al., 2014; Yu et al., 2023; Yang et al., 2021). Given historical observations, it is valuable if we can know the future values ahead of time. Correspondingly, TSF has been firmly established as an active research field, witnessing the proposal of numerous methods.
Figure 1. Visualization of data with different characteristics.
Time series organize data points chronologically and are either univariate or multivariate depending on the number of variables in each data point. Accordingly, TSF methods can be classified as either Univariate Time Series Forecasting (UTSF) or Multivariate Time Series Forecasting (MTSF) methods. Among early methods, Autoregressive Integrated Moving Average (ARIMA) (Box and Pierce, 1970) and Vector Autoregression (VAR) (Toda and Phillips, 1994) are arguably the most popular univariate and multivariate forecasting methods, respectively. Subsequent methods that exploit machine learning, e.g., XGBoost (Chen and Guestrin, 2016; Zhang et al., 2021) and Random Forest (Breiman, 2001; Mei et al., 2014) offer better performance than the early methods. Most recently, methods based on deep learning have demonstrated state-of-the-art (SOTA) forecasting performance on a variety of datasets (Zhou et al., 2021; Nie et al., 2022; Wu et al., 2022; Zhang and Yan, 2022; Chen et al., 2024; Yao et al., 2023; Wu et al., 2021b, 2023; Zhao et al., 2023; Pedersen et al., 2020; Cheng et al., 2023; Cirstea et al., 2022b, 2019, 2021; Lin et al., 2023a; Miao et al., 2024).
@@
Issue 1. Insufficient Coverage of Data Domains. Time series from different domains may exhibit diverse characteristics. Figure 1a depicts a time series from the environment domain called AQShunyi (Zhang et al., 2017) that records temperature information at hourly intervals, exhibiting a distinct seasonal pattern. This pattern is reasonable in this scenario because temperatures in nature often cycle around the year. Figure 1b shows a time series from FRED-MD (McCracken and Ng, 2016) belongs to economic domain that describes the monthly macroeconomic from 114 regional, national, and international sources with a clear increasing tendency. This may be attributed to overall economic stability with minimal fluctuations, reflecting sustained growth in the macroeconomic indicators. Figure 1c depicts a series among Electricity (Trindade, 2015) which comes from electricity domain and has a significant change in the data at a certain point in time, which might indicate an abrupt event, etc. However, these simple patterns are only the tip of the iceberg, and time series from different domains may exhibit much more complex patterns that either combine the above characteristics or are entirely different. Therefore, using only limited domains results in limited coverage of time series characteristics, which cannot offer a full picture.
@@
Datasets VAR LR PatchTST NLinear FEDformer Crossformer
NASDAQ 0.462 0.616 0.567 0.522 0.547 0.745
Wind 0.620 0.583 0.652 0.640 0.697 0.590
ILI 1.012 4.856 0.835 0.919 1.020 1.096
@@
Table 2. Impact of batch sizes with “drop last.”
Size PatchTST DLinear FEDformer
1 0.4203 0.4874 0.4120
32 0.4138 0.4831 0.4084
64 0.3999 0.4726 0.4022
128 0.3750 0.4539 0.3921
256 0.3561 0.4360 0.3825
512 0.3483 0.4251 0.3736
Issue 3. Lack of consistent and flexible pipelines. The performance of different methods varies with changing experimental settings, e.g., splits among training/validation/testing data, the choice of normalization method, and the selection of hyper-parameter settings. For example, implementations of existing methods often employ a “Drop Last” trick in the testing phase (Zhou et al., 2022b; Nie et al., 2022; Wang et al., 2022). To accelerate the testing, it is common to split the data into batches. However, if we discard the last incomplete batch with fewer instances than the batch size, this causes unfair comparisons. For example, in Figure 4, the ETTh2 (Zhou et al., 2021) has a testing series of length 2,880, and we need to predict 336 future time steps using a look-back window of size 512. If we select the batch sizes to be 32, 64, and 128, the number of samples in the last batch are 17, 49, and 113, respectively. Discarding those last-batch testing samples is inappropriate unless all methods use the same strategy. Table 4 shows the testing results of PatchTST (Nie et al., 2022), DLinear (Zeng et al., 2023) and FEDformer (Zhou et al., 2022b) on the ETTh2 with different batch sizes and the “Drop Last” trick turned on. We observe that the performance of the methods changes when varying the batch size. In addition, the pipelines in most benchmarks are inflexible and do not support simultaneous evaluation of statistical learning, machine learning, and deep learning methods. We argue that it is crucial to ensure a consistent and flexible pipeline so that methods are evaluated in the same setting, thereby improving the fairness of the findings.
Robust and extensive benchmarks enable researchers to evaluate proposals for new methods more rigorously across diverse range of datasets, which is crucial for advancing the state-of-the-art (Tan et al., 2020; Qiao et al., 2024). For example, ImageNet (Deng et al., 2009), that encompasses a substantial dataset has been instrumental in ensuring progress in computer vision. ImageNet has established itself as a standard for assessing methods in image processing due to its support for rigorous evaluation. Table 3 compares existing benchmarks for TSF according to seven properties. No existing benchmark possesses all properties.
@@
Benchmark Property Univariate forecasting Multivariate forecasting Statistical method Machine learning method Deep learning method Taxonomy of data Flexible & scalable pipeline
M3 (Makridakis and Hibon, 2000) $\surd$
$\times$
$\surd$ $\surd$
$\times$
$\times$
$\times$
M4 (Makridakis et al., 2018) $\surd$
$\times$
$\surd$ $\surd$ $\surd$
$\times$
$\times$
LTSF-Linear (Zeng et al., 2023)
@@
$\surd$
@@
[…]
TSlib (Wu et al., 2022) $\surd$ $\surd$
@@
[…]
BasicTS (Liang et al., 2022)
@@
$\surd$ $\surd$
@@
Libra (Bauer et al., 2021) $\surd$
$\times$
$\surd$ $\surd$
$\times$
$\times$
[…]
TFB (Ours) $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$ $\surd$
@@ 3. Preliminaries
(4) $\footnotesize F=\mathit{Catch22}(X)$
@@
(5) $\footnotesize P=\left\{r(F^{i},F^{j})\mid 1\leq i\leq N,i+1\leq j\leq N,i,j\in N^{\ast}\right\}$
@@ 4. TFB: BENCHMARK DETAILS
  • 1
    $|TS|\textless 300$: the length of univariate time series <300
  • 2
    $F$: the forecasting horizon
@@ 4.1.1. Dataset overview
Univariate time series. The univariate datasets are carefully curated from 16 open-source datasets, thus covering dozens of domains. To fully reflect the complexity of real-world time series, we employed Principal Feature Analysis (PFA) (Lu et al., 2007), a variation of Principal Component Analysis (PCA) (Bro and Smilde, 2014). PFA preserves the original values of individual time series data points. We employ the concept of explained variance, representing the ratio between the variance of a single time series and the sum of variances across all individual time series. A threshold $t$ for the explained variance is set to 0.9. This implies that for each dataset, we choose to retain the minimum number of time series required to encompass 90% of the variance contributed by the remaining time series. As a result, the selected data exhibits pronounced heterogeneity. Compared to datasets with strong homogeneity, it can better reflect methods performance. In the end, we select 8,068 time series to enable the combined dataset to capture the diversity of real-world time series. Statistical information is reported in Table 4.
@@ 4.2. Comparison Methods
To investigate the strengths and limitations of different forecasting methods, we include 22 methods that can be categorized as statistical learning, machine learning, and deep learning methods. In terms of statistical learning methods, we include ARIMA (Box and Pierce, 1970), ETS (Hyndman et al., 2008), Kalman Filter (KF) (Harvey, 1990), and VAR (Toda and Phillips, 1994). Among the machine learning methods, we include XGBModel (XGB) (Herzen et al., 2022), LinearRegression (LR) (Kedem and Fokianos, 2005; Herzen et al., 2022), and Random Forest (RF) (Breiman, 2001). Finally, we further split the deep learning methods into RNN-based models (RNN (Kim et al., 2021)), CNN-based models (MICN (Wang et al., 2022), TimesNet (Wu et al., 2022), and TCN (Bai et al., 2018)), MLP-based models (NLinear (Zeng et al., 2023), DLinear (Zeng et al., 2023), TiDE (Das et al., 2023), N-HiTS (Challu et al., 2023), and N-BEATS (Oreshkin et al., 2019)), Transformer-based models (PatchTST (Nie et al., 2022), Crossformer (Zhang and Yan, 2022), and FEDformer (Zhou et al., 2022b), Non-stationary Transformer (Stationary) (Liu et al., 2022), Informer (Zhou et al., 2021), and Triformer (Cirstea et al., 2022a)), and Model-Agnostic models (FiLM (Zhou et al., 2022a)).
@@ 4.3.1. Evaluation strategies
Rolling forecasting. In Rolling forecasting, illustrated in Figure 6(b), the blue squares represent historical data, the green squares represent the forecasting horizon, and the white squares represent the remaining data in the time series. During rolling forecast, excluding the last iteration, the historical data expands by a fixed number of steps (called the stride) in each iteration. In the final iteration, the historical data is extended to cover the entire time series along with the forecasting horizon. In each iteration of the inference process, the forecasting model is applied based on the historical data to forecast the designated forecasting horizon. Then, we calculate the average of the error metrics for each iteration. In statistical learning methods (e.g., ARIMA (Box and Pierce, 1970), ETS (Federico Garza, 2022)), it is common to use the entire or a subset of the historical data for training and then make forecasting during each iteration. In contrast, in deep learning or machine learning methods, each iteration often involves using only the last portion of the historical data, with a length equal to the look-back windows size, for inference to make forecasting. The current evaluation strategy for TSF, such as TimesNet (Wu et al., 2022), aligns with our defined criteria.